Circular No. 6109

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) Phone 617-495-7244/7440/7444 TWX 710-320-6842 ASTROGRAM CAM

MARSDEN@CFA.HARVARD.EDU or GREEN@CFA.HARVARD.EDU (science)

LEONID METEORS 1994 Reports of enhanced activity of the Leonid meteors on Nov. 17–18,

both via visual observation (despite the full moon) and radio monitoring, have been received from P. Jenniskens (Ames Research Center, NASA; and Dutch Meteor Society), K. Suzuki (Aichi, Japan) and T. Nakamura (Radio Atmospheric Science Center, Kyoto University); and P. Brown (University of Western Ontario). Both the true rates and the peak time are difficult to ascertain, though it seems possible that the zenithal hourly rate was ∼ 100 around Nov. 18.2 UT ± 0.4 day (corresponding to solar longitude ∼ 235^{o}.7 ± 0^{o}.4, equinox 2000.0).

CH CYGNI A. Skopal, Astronomical Institute, Skalnaté Pleso Observatory; and

H. M. Lloyd and M. F. Bode, John Moores University, Liverpool, write: “Skalnaté Pleso photometry and International Ultraviolet Explorer (IUE) observations show an eclipse of the active component by the cool component in this symbiotic system. Photometric observations indicate a deep minimum of ∼ 3 mag in the U band: July 29.86 UT, U = 6.562, B = 7.620, V = 7.279; Aug. 29.10, 6.815, 7.936, 7.583; Oct. 1.02, 7.605, 8.364, 7.731; Oct. 12.87, 8.966, 9.401, 8.324; Oct. 15.80, 9.667,9.942, 8.589; Oct. 27.88, 9.136, 9.469, 8.440; Nov. 9.83, 9.322 9.575, 8.536; Nov. 23.72, 9.268, 9.764, 8.682. An IUE spectrum obtained on Oct. 14 is similar in both the shape and level of the continuum to that obtained on 1992 Nov. 24, during what would have been the last eclipse of a 756-day-period system. However, the latest spectrum shows strong emission in the O I 130.5-nm line, which was absent on 1992 Nov. 24. The ultraviolet continuum has decreased relative to 1993 May 28 (IAUC 5815) by a factor of ∼ 4 between 300 and 330 nm, to 5 × 10^{−13} erg cm^{−2} s^{−1} A^{−1}, and by a factor of two around 150 nm, to 1.2 × 10^{−12} erg cm^{−2} s^{−1} A^{−1}. We estimate that the end of the eclipse should occur in December. Positions of the latest minima (1992 Oct. and 1994 Oct.–Nov.) agree well with times of the spectroscopic conjuction of the symbiotic pair in the proposed triple-star model of CH Cyg.”

SUPERNOVA 1994ae IN NGC 3370 S. Nakano, Sumoto, Japan, forwards the following magnitude esti-

mates: Nov. 29.740 UT, 12.5 (K. Takamizawa, Saku-machi, Nagano-ken; photographic); 29.836, 12.5 (R. Kushida, Yatsugatake South Base Observatory; unfiltered CCD); 30.854, 12 (K. Ikeya, Shizuoka; visual).

1994 December 1 Daniel W. E. Green